OpenStudy (anonymous):

Find the limit: lim as t → infinity (sqr t + t^2) / ( 5t − t^2)

OpenStudy (anonymous):

Divide everything by \(t^2\) since if is the highest degree term in the denominator.

OpenStudy (anonymous):

use l'hopitals rule

OpenStudy (anonymous):

is it \[\frac{\sqrt{t+t^2}}{5t-t^2}\] or is it \[\frac{\sqrt{t}+t^2}{5t-t^2}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

then use nothing it is the ratio of the leading coefficients

OpenStudy (anonymous):

the first time you do the limit it goes to the form inf/inf so use l'hopitals

OpenStudy (anonymous):

I get ( sqr t ) / (5t) is that the answer?

OpenStudy (anonymous):

oh heck no you don't use any calculus for this

OpenStudy (anonymous):

No calculus for limits?

OpenStudy (anonymous):

and also @latona heck no it is the ratio of the leading coefficients

OpenStudy (anonymous):

this is what i got when I divided by t^2

OpenStudy (anonymous):

no, not for limits as \(x\) goes to infinity, which is a synonym for horizontal asymptotes

OpenStudy (anonymous):

the leading coefficient of the numerator is 1, the leading coefficient of the denominator is -1, and the ration is \(\frac{1}{-1}=-1\)

OpenStudy (anonymous):

*ratio

OpenStudy (anonymous):

Kind of confused now how to approach the problem. Can you help further, anyone? Thanks.

OpenStudy (anonymous):

do you know what l'hopitals rule is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know how to divide?

OpenStudy (anonymous):

Divide everything by \(t^2\) like I said.

OpenStudy (anonymous):

when I divide everything by t^2, I get \[\frac{ \sqrt{t} }{ 5t }\]

OpenStudy (anonymous):

the numerator has \(t^2\) and the denominator as \(-t^2\) the other stuff is unimportant since the degrees are all smaller than 2 \[\frac{t^2}{-t^2}=-1\] is your asymptote